OpenStudy (anonymous):

Use the linear approximation (1+x)^kapprox 1+kx to find an approximation for the function f(x) for values of x near zero. a.) f(x)=(1-x)^(8) approx b.) f(x)=-8/(1-x) = approx c.) f(x)= 1/(sqrt(1+x)) = approx d.) f(x)=sqrt(4+x^2) approx = e.) f(x)=(6+3x)^(1/3) approx = How do I go about attacking a problem like this?

OpenStudy (anonymous):

Use the linear approximation (1+x)^k approx = 1+kx to find an approximation for the function f(x) for values of x near zero.

OpenStudy (inkyvoyd):

for the first one it should be straightfoward.... just use (-x) in palce of x for the approximation given and your k value is 8

OpenStudy (anonymous):

Ohh hahah I think I'm making this more difficult than it is

OpenStudy (inkyvoyd):

okay. for B note that 1/y=y^(-1)

OpenStudy (anonymous):

What would b =? I know all the derivative rules and everything I just don't really get this for some reason hahah.

OpenStudy (inkyvoyd):

b=-8*(1-x)^-(1)

OpenStudy (anonymous):

So, -8*(1+x)! Awesome I think I get it. Thank ya

OpenStudy (anonymous):

For D, I know that (4+x^2)^(1/2) so (4+(1/2)x^2). But that's wrong. Can you catch my error? I'm sorry I know this shouldn't be this difficult!

OpenStudy (anonymous):

take 4 out

OpenStudy (anonymous):

\(\sqrt{4+x^2}=2\sqrt{1+(x/2)^2}\)